Express 45% and 78% as a fraction.Express 45% and 78% as a fraction.Express 45% and 78% as a fraction.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

45% = 45/100 = 9/20

78% = 78/100 = 39/50

Thanks